# whose only singularity is a double point

### 04 Mar 2026 v1

@@ Abstract.
The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ that have a unique singular point, which has multiplicity 2; see Example 3.
@@
For $d\geq 10$ the curves $C_{d}$ cannot be lifted to characteristic 0 in a way that preserves the standard resolution-of-singularities blow-up sequence; see Remark 7. This is inconsistent with my understanding of some steps of the program outlined in [Ish25]; see Remark 7. However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic; see Remark 8. This question is one of the main topics of [Ish25].
@@
Fix a prime $p$, let $q$ be a $p$-power and $r$ a multiple of $p$. Let $C^{\circ}_{q,r}\subset{\mathbb{A}}^{2}$ be the the image of the map
$\phi_{q,r}:t\mapsto\bigl(t^{q},t^{q+r}+t\bigr)=:(x,y),$
and $C_{q,r}\subset{\mathbb{P}}^{2}$ its closure.
@@
For the curves $C_{q,r}$ in Example 1 the following hold.
@@
    $\phi_{q,r}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
@@
    $C^{\circ}_{q,r}$ is a smooth, rational curve, and
@@
    $C_{q,r}$ has a single point at infinity, which has multiplicity $r$.
Proof. The derivatve of $\phi_{q,r}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,r}$ is obtained using
(2.4) $y^{q}=t^{q(q+r)}+t^{q}=x^{q+r}+x.$
The unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
(2.5) $z^{r}=x^{q+r}+xz^{q+r-1}.\qed$
@@
It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{15}{28}(\deg B)^{2}$ [GZN00].
@@
For $d\leq 6$ [Yos79] gives degree $d$, rational, plane curves $Y_{d}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with a single singular point, which has multiplicity 2.
For $d=6$ the equation is explicit, but rather complicated. The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible.
@@
The smallest counterexample is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity. By contrast, any reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{60}$ singularity by Lemma 4.
@@
However, the log canonical threshold of the surface $X(g)\subset{\mathbb{A}}^{3}$ constructed from $C_{2^{n},2}$ is $3/(2^{n}+2)$. It is computed by the exceptional divisor of the first blow-up, and a general point of the same multiplicity has the same log canonical threshold.
Note, however, that these non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways. The next examples show that liftability may frequently hold.
@@ Acknowledgments.
I thank M. Schütt and I. Shimada for numerous e-mails explaining supersingular K3 surfaces. Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).

### 16 Mar 2026 v2

@@ Abstract.
The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ or $d=2^{n}$ that have a unique singular point, which has multiplicity 2; see Example 3.
@@
For $d\geq 8$ the curves $C_{d}$ cannot be lifted to characteristic 0 in a way that preserves the standard resolution-of-singularities blow-up sequence; see Remark 7. This is inconsistent with my understanding of some steps of the program outlined in [Ish25]; see Remark 7. However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic, which is one of the main topics of [Ish25]; see Remark 8.
@@
Fix a characteristic $p>0$. Let $q$ be a $p$-power, and $g(u)\in{\mathbb{F}}_{p}[u]$ a polynomial. Let $C^{\circ}_{q,g}\subset{\mathbb{A}}^{2}$ be the the image of the map
$\phi_{q,g}:t\mapsto\bigl(t^{q},g(t^{p})+t\bigr)=:(x,y),$
and $C_{q,g}\subset{\mathbb{P}}^{2}$ its closure.
@@
For the curves $C_{q,g}$ in Example 1 the following hold.
@@
    $\phi_{q,g}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
@@
    $C^{\circ}_{q,g}$ is a smooth, rational curve, and
@@
    $C_{q,g}$ has a single point at infinity. If $r:=p\cdot\deg g-q$ is positive, it has multiplicity $r$.
Proof. The derivatve of $\phi_{q,g}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,g}$ is obtained using
(2.4) $y^{q}=g(t^{qp})+t^{q}=g(x^{p})+x.$
If $r:=p\cdot\deg g-q$ is positive, then the unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
(2.5) $z^{r}=g(x^{p})+xz^{q+r-1}.\qed$
@@
Setting $p=2,q=2^{n}$ and $g(u)=u^{2^{n-1}-1}$, we get
$C_{2^{n}}:=\bigl(y^{2^{n}}=z^{2}+z^{2^{n}-1}\bigr)\subset{\mathbb{P}}^{2}.$
It is a rational curve of degree $2^{n}$. It has a unique singular point of type $A_{m}$, where $m=(2^{n}-1)(2^{n}-2)$.
@@
It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{7}{12}(\deg B)^{2}$ [Ore12, Sec.5].
@@
For $d\leq 6$ [Yos79] gives degree $d$, rational, plane curves $Y_{d}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with a single singular point, which has multiplicity 2. The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible. Simpler equations are given in [ABCRCA02, 5.8].
@@
The smallest counterexample is $C_{8}$, which has degree $8$ and an $A_{42}$ singularity. A reduced, octic curve $B_{8}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{37}$ singularity by Lemma 4.
The next one is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity. A reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{61}$ singularity by Lemma 4.
@@
The program of Ishii aims to understand log canonical thresholds in positive characteristic [Ish25].
@@
However, the log canonical threshold of the surface $X(g)\subset{\mathbb{A}}^{3}$ constructed from $C_{2^{n},2}$ is $3/(2^{n}+2)$. It is computed by the exceptional divisor of the first blow-up, and a general singular point with the same multiplicity has the same log canonical threshold. Thus these curves do not seem to give further information about the program in [Ish25].
Note, however, that the above non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways. The next examples show that liftability may frequently hold.
@@
The curves $C_{2^{n},2}$ also lead to supersingular double planes.
10Supersingular double planes.
Over ${\mathbb{F}}_{2}$ consider the surfaces
(10.1) $S_{r}:=\bigl(u^{2}+u(zy^{2r}+x^{2r+1})+xz^{4r+1}+z^{2}y^{4r}+x^{4r+2}=0\bigr)\subset{\mathbb{P}}(1,1,1,2r+1).$
In the $y=1$ chart, the equation is
(10.2) $g:=u^{2}+u(z+x^{2r+1})+xz^{4r+1}+z^{2}+x^{4r+2}=0.$
Thus $g_{u}=z+x^{2r+1},\ g_{z}=u+xz^{4r},\ g_{x}=ux^{2r}+z^{4r+1}.$ The first 2 define the monomial curve $M_{r}$ which is the image of
$\phi_{r}:t\mapsto\bigl(t,t^{2r+1},t^{8r^{2}+4r+1}\bigr),$
and $g_{x}$ vanishes identically along $M_{r}$. Thus
$\begin{array}[]{rcl}{\mathbb{F}}_{2}[x,z,u]/(g,g_{x},g_{z},g_{u})&\cong&{\mathbb{F}}_{2}[t]/\bigl(g(t,t^{2r+1},t^{8r^{2}+4r+1})\bigr)\\ &\cong&{\mathbb{F}}_{2}[t]/\bigl(t^{8r^{2}+6r+2}(t^{8r^{2}+2r}+1)\bigr).\end{array}$
Thus $S_{r}$ has an $A_{8r^{2}+6r+1}$ singularity at $t=0$. For the other singularities, write $2r=qr^{\prime}$ where $q$ is a $2$-power and $r^{\prime}$ is odd. Then we have $A_{q-1}$ singularities when $t$ is a $(4rr^{\prime}+r^{\prime})^{th}$ root of unity. There are no other singular points.
If $r$ is a $2$-power, then the minimal resolution $\bar{S}_{r}\to S_{r}$ contains $8r^{2}+6r+1+(4r+1)(2r-1)=16r^{2}+4r$ exceptional curves.
Over the line $(z=0)$ the equation (10.2) becomes $u^{2}+ux^{2r+1}+x^{4r+2}=0$. Setting $v:=ux^{-2r-1}$ we get $v^{2}+v+1=0$. If $\alpha_{1},\alpha_{2}$ are its roots, then the preimage of $(z=0)$ has 2 components
$B_{i}:=\bigl(u+\alpha_{i}x^{2r+1}=z=0\bigr).$
Thus the Picard number of $\bar{S}_{r}$ is at least $16r^{2}+4r+2$.
This equals the second Betti bumber of a double plane ramified along a smooth curve of degree $4r+2$. Thus the surfaces $\bar{S}_{r}$ are supersingular for every $r=2^{n}$, with Picard number $16r^{2}+4r+2$.
Example 11.
The surface $S_{1}$ is a K3 surface with $A_{15}+5A_{1}$ singularities. Resolving the $A_{15}$ singularity, the birational transforms of the $B_{i}$ become smooth rational curves, and the dual configuration is
$\begin{array}[]{ccccccccccccc}&&&&\bar{B}_{1}&&&&\bar{B}_{2}&&&&\\ &&&&|&&&&|&&&&\\ C_{1}&-&C_{2}&-&C_{3}&-&\cdots&-&C_{13}&-&C_{14}&-&C_{15}\end{array}$
From this we compute that $(B_{1}\cdot B_{2})=\frac{9}{16}$ and $(B_{1}^{2})=(B_{2}^{2})=\frac{7}{16}$.
Contracting all but the curve $C_{8}$ gives a supersingular K3 surface $S^{\prime}_{1}$ with singularities $2E_{8}+5A_{1}$, and Picard number $1$. The image of $C_{8}$ is a rational curve $C^{\prime}_{8}$ with 2 cusps at the $E_{8}$-points, and $(C^{\prime}_{8}\cdot C^{\prime}_{8})=2$. Thus $S^{\prime}_{1}$ is a K3 surface of degree 2. This is the 4th surface on the list of [Shi04] in characteristic 2.
@@ Acknowledgments.
I thank M. Schütt and I. Shimada for numerous e-mails explaining supersingular K3 surfaces, and S. Orevkov and M. Zaidenberg for references. Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).